OpenStudy (anonymous):

A car traveling at 60 km/h will skid 30 m when its brakes are locked. If the same car is traveling at 180 km/h, what will be its skidding distance? can you please Show how you get the aswer

OpenStudy (chmvijay):

http://answers.yahoo.com/question/index?qid=20081214034303AAeJxx9 this will help you :)

OpenStudy (chmvijay):

LOL did it help or shall i explain it :)

OpenStudy (anonymous):

can you explain it pls

OpenStudy (mos1635):

I propose energy-work theorem:\[\frac{ 1 }{ 2 }*M*V _{1}^2 = T(friction)* d _{2}\]

OpenStudy (mos1635):

\[\frac{ 1 }{ 2 }* M *V _{2}^2=T(friction)∗d2\]

OpenStudy (mos1635):

divide

OpenStudy (mos1635):

...\[\left( \frac{ V _{1} }{ V _{2} } \right)^2 =\frac{ d _{1} }{ d _{2} }\]

OpenStudy (mos1635):

.........

OpenStudy (anoop27):